Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
5x + 35y = -10
15x + 5y = 170

Answer by mananth(16946)

5 x + 35 y = -10 .............1
15 x + 5 y = 170 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -7
5 x 35 y = -10
-105 x -35 y = -1190
Add the two equations
-100 x = -1200
/ -100
x = 12
plug value of x in (1)
5 x + 35 y = -10
60 + 35 y = -10
35 y = -10-60
35 y = -70
y = -2